Find an equation of the line having the given slope and containing the given point. Express your answer in the form y + b or y = mx + b.
m = 0, (0, -7).
The equation of the line is y = _?

Answer by ewatrrr(24785)


Hi
Finding an equation of the line with m = 0 and y-intercept (0-7)
Using the standard slope-intercept form for an equation of a line y = mx + b
where m is the slope and b the y-intercept. m = 0 and b = -7
y = 0*x - 7
y = -7 equation of the line in the y = mx+b format